OpenStudy (anonymous):

What is the value of z so that -9 and 9 are both solutions of x^2+z=103? A)-22 B)3 C)22 D)184

OpenStudy (anonymous):

x^2=103-z Soive 103-z = 81 z = 22

OpenStudy (anonymous):

x^2 + 22 = 103 x^2 = 81 x = +/- 9